OpenStudy (anonymous):

write an equation of the line that passes through point p and is parallel to the line with the given equation P(0,-1), y=-2x+3 if you plz could help me figure out this problem and explain it so i can do the rest of them?

OpenStudy (ranga):

y = -2x + 3 Compare this to y = mx + b where m is the slope. The slope of y = -2x + 3 is -2. Parallel lines all have the same slope. Therefore, a line parallel to y = -2x + 3 will also have a slope of -2. This parallel line passes through (0,-1). The point-slope form of a line is: y - y1 = m(x - x1) where (x1, y1) is the point it passes through and m is the slope. Here (x1,y1) is (0,-1) and m = -2. Substitute in y - y1 = m(x - x1) and simplify.